# RePrompt: Planning by Automatic Prompt Engineering for Large Language Models Agents

### v1

@@ Appendix A Pseudo Code for RePrompt
1:function Train
2:     $Prompt\leftarrow InitialPrompt.$
3:     for batch, x from Dataloader do
4:         $Response\leftarrow Act(Prompt,x)$
5:         $Loss\leftarrow Summarize(Response,x,Prompt)$
6:         $Prompt\leftarrow Optimizer(Loss,Prompt)$
@@
Appendix B Prompts for RePrompt
@@
One example of the placeholder is: <̈Original Prompt Start>.̈ You need to replace this placeholder with the original prompt.
@@
4. Finish your output with your final prompt, in the format of: "Based on the above analysis, the improved prompt is:
n".
@@
- If it is already a thought, then a check on whether the thought is still workable in the given scenario is very helpful. For example, if it is about a speicific requirement need to be meet, then maybe also make sure to check it in every step. However, make sure this does not limit what the feedback can provide, and using words like s̈pecificallyẗo remind such a check.
@@
Please provide a detailed and complete response without omitting any information or use "…" or "[…]"to replace any part of the prompt. Again, ensure that no information is omitted or summarized.
@@
Here, we provide all the prompts used in our paper. The prompt used to summarize the loss is provided in Fig. 3, the prompt used to optimize the prompt is provided in Fig. 5, and the prompt used to replace the placeholders that could be accidently included, which is discussed in Sec. 5.1, is provided in Fig. 4.
Appendix C Optimized Prompt from RePrompt
@@
Figure 7: The prompt of TravelPlanner optimized after 1 epoch.The main changes are highlighted in blue.
@@
You must use Finish to indict you have finished the task. And each action only calls one function once. Given information: text Query: queryscratchpad
Figure 8: The prompt of TravelPlanner optimized after 5 epochs.
Here, to help reproducibility, we provide all the optimized prompts that leads to the results shown in Table. 1 and Table. 2. It needs to be addressed that even if the prompt, the temperature, the model used, and the seed are the same, OpenAI APIs still do not guarantee that the generated output will be exactly the same every time. This will greatly affect the final results in our case, given that Reflexion [Shinn et al., 2023] is used to provide the feedback, and a small change on the earlier reflection can lead to completely different results in the end. Due to unknown reasons, we found that the reflection module in Reflexion [Shinn et al., 2023] can provide completely useless and even wrong suggestions to the LLM and lead to very bad results. If this happens, we suggest rerunning the baseline model without RePrompt to make sure the OpenAI is providing correct feedbacks. However, the final conclusion, especially the relatively superior of our model, is always found to be true in our experiments.
In Fig. 6, we provide the optimized prompt for PDDL action generation. In Fig. 7 and Fig. 8, we provide the optimized prompt for the TravelPlanner environment [Xie et al., 2024].

### v2

@@ Appendix C Pseudo Code for RePrompt
1:  function Train()
2:  $Prompt\leftarrow\text{Initial Prompt}$
3:  for (batch, x) in Dataloader do
4:     $Response\leftarrow\textsc{Act}(Prompt,x)$
5:     $Loss\leftarrow\textsc{Summarize}(Response,x,Prompt)$
6:     $Prompt\leftarrow\textsc{Optimizer}(Loss,Prompt)$
@@
Appendix D Prompts for RePrompt
{mdframed}
[…]
==== Given information: {information}
Query: {query}
The generated plan is: {plan}
Figure 4: The checker prompt for PromptAgent. The prompt should be fed into an LLM, which in this paper is GPT4-turbo, to get the judgement of whether the current result is correct or wrong.
@@ Appendix D Prompts for RePrompt
One example of the placeholder is: “ $\langle$ Original Prompt Start $\rangle$ ”. You need to replace this placeholder with the original prompt.
@@
4. Finish your output with your final prompt, in the format of: "Based on the above analysis, the improved prompt is: ".
@@
- If it is already a thought, then a check on whether the thought is still workable in the given scenario is very helpful. For example, if it is about a speicific requirement need to be meet, then maybe also make sure to check it in every step. However, make sure this does not limit what the feedback can provide, and using words like “specifically" to remind such a check.
@@
Please provide a detailed and complete response without omitting any information or use “…" or “[…]"to replace any part of the prompt. Again, ensure that no information is omitted or summarized.
@@
Here, we provide all the prompts used in our paper. The prompt used to summarize the loss is provided in Fig. 5, the prompt used to optimize the prompt is provided in Fig. 7, and the prompt used to replace the placeholders that could be accidently included, which is discussed in Sec. 5, is provided in Fig. 6.
Appendix E Optimized Prompt from RePrompt
Here, to help reproducibility, we provide all the optimized prompts that leads to the results shown in Table. 1 and Table. 2. It needs to be addressed that even if the prompt, the temperature, the model used, and the seed are the same, OpenAI APIs still do not guarantee that the generated output will be exactly the same every time. This will greatly affect the final results in our case, given that Reflexion [Shinn et al., 2023] is used to provide the feedback, and a small change on the earlier reflection can lead to completely different results in the end. Due to unknown reasons, we found that the reflection module in Reflexion [Shinn et al., 2023] can provide completely useless and even wrong suggestions to the LLM and lead to very bad results. If this happens, we suggest rerunning the baseline model without RePrompt to make sure the OpenAI is providing correct feedbacks. However, the final conclusion, especially the relatively superior of our model, is always found to be true in our experiments. And for a fair comparison and to match the results from the TravelPlanner paper, all the results reported in the paper are the best among the 3 trials over time (Best-of-3).
In Fig. 8, we provide the optimized prompt for PDDL action generation. In Fig. 9 and Fig. 11, we provide the optimized prompt for the TravelPlanner environment [Xie et al., 2024]. In Fig. 10, we provide the optimized prompt for Meeting Planning task. In Fig. 12, we provide the optimized prompt generated by PromptAgent Wang et al. [2024] for TravelPlanner environment.
@@ Appendix E Optimized Prompt from RePrompt
Figure 9: The prompt of TravelPlanner optimized after 1 epoch of Reprompt. The main changes are highlighted in blue.
{mdframed}
[linecolor=black, linewidth=1pt] You are visiting San Francisco for the day and want to meet as many friends as possible. Solve the problem by considering various different schedules and picking the best one to optimize your goals.
1. **Identify mandatory time constraints**: Start by listing all fixed commitments (e.g., travel to/from the city, meal breaks, pre-scheduled events). Calculate total time consumed by these. 2. **Maximize flexible availability**: Subtract mandatory time from your total available hours. This remaining time will be used for friend meetings. 3. **Evaluate scheduling options**: a. **Identify friend clusters**: - **First analyze proximity efficiency factors**: • Evaluate mandatory commitments’ locations **and time windows** • Identify transit corridors/routes connecting multiple commitments • Map time-sensitive opportunities (e.g., friends available during/after nearby commitments) - Form initial clusters by: • Grouping friends within 15-minute transit of **any mandatory commitment location or transit corridor** • Prioritizing clusters that align with commitment time windows (e.g., friends near lunch spot during meal break) - Then map remaining friends into geographic clusters based on density - Calculate transit time between all clusters and to/from mandatory commitments b. Prioritize clusters that: - **Use a scoring system**: • Calculate cluster scores by combining: - **Proximity score**: - 8 points if $\leq$10min transit to nearest commitment/corridor - 5 points if 11-20min - 2 points if 21-30min - 0 points if >30min - **Density score**: (Number of friends in cluster) × (3 ÷ cluster radius in miles) - Example: 6 friends in 0.6-mile radius = 6 × (3/0.6) = 30 points • Prioritize clusters with the highest combined scores first • **Re-evaluate scores after completing each mandatory commitment** to account for new proximity opportunities - **Structure clusters to**: • Start with clusters **adjacent to early mandatory commitments** to establish proximity early • **After completing any mandatory commitment, remain in its cluster** to meet nearby friends and avoid backtracking • Sequence clusters **along transit corridors in one direction** (e.g., north-to-south) to minimize crisscrossing • Place clusters scoring $\geq$35 total points **immediately after mandatory commitments** when time windows allow extended stays - Enable efficient sequential visitation with minimal internal transit time c. Compare scenarios where you: - Start early vs. late - Sequence clusters by proximity to mandatory commitments first - Adjust meeting durations within clusters 4. **Select the optimal schedule**: Choose the option that maximizes the number of friends met while respecting all constraints.
query
Your response should start with ’SOLUTION:’.
Figure 10: The prompt of Meeting Planning optimized after 5 epoch of Reprompt. The main changes are highlighted in blue.
@@
Given information: {text}
Query: {query}{scratchpad}
Figure 11: The prompt of TravelPlanner optimized after 5 epochs of Reprompt. The main changes are highlighted in blue.
{mdframed}
[linecolor=black, linewidth=1pt] You are a proficient planner with a keen eye for detail and practicality. Your task is to create a comprehensive travel plan that adheres to the provided budget and timeframe, ensuring a diverse and enjoyable experience. The plan should include specific flight numbers, restaurant names, hotel names, and attraction details, all of which must be derived from the provided data. Follow the format shown in the example, and ensure that all details are sensible and feasible.
When planning, consider the following guidelines:
- Ensure meal diversity by not repeating restaurant choices for different meals.
- Select transportation options that are practical and feasible, considering the distance and geography.
- Provide complete information, including all meals for each day and any in-city transportation if necessary.
- Adhere to the budget, allocating funds across flights, accommodations, meals, and attractions.
- Check for any minimum stay requirements or house rules for accommodations.
- Use the symbol ’-’ to indicate when information is unnecessary, such as after returning to the departure city or when no transportation is needed within the current city.
Your planning process should consist of alternating Thought, Action, and Observation steps:
- Thought: Reason about the current situation and what needs to be planned next.
- Action: Perform one of two types of actions: (1) CostEnquiry[Sub Plan]: Calculate the cost of a detailed sub-plan for a complete one-day plan. Input the number of people and the plan in JSON format.
(2) Finish[Final Plan]: Indicate the completion of the task by submitting a final, complete plan as an argument.
- Observation: Reflect on the information received from the actions and adjust the plan accordingly.
Remember, each action should only call one function once, and you must use Finish to indicate you have finished the task.
Here is an example for reference:
***** Example *****
{Example content}
***** Example Ends *****
Now, let’s begin planning based on the given information and query. Keep in mind that the plan should be logical, feasible, and within the specified constraints. Good luck!
{reflections}
Given information: {text} Query: {query}{scratchpad}
Figure 12: The prompt of TravelPlanner optimized by PromptAgent. Unlike Reprompt, majority prompt has been changed and thus we do not do further highlight.